# Exploring Cultural Variations in Moral Judgments with Large Language Models

### v1

@@ 5 Discussion and Conclusion
Our findings show that language models vary considerably in how well they replicate cross-cultural moral judgments, as captured in the WVS and PEW surveys. Larger or instruction-tuned models, such as Falcon-40I, Gemma-9, and GPT4o, frequently demonstrate higher correlations with aggregated human survey responses. In contrast, some models, including Qwen-0.5 and Llama2-70, yield systematically negative correlations, suggesting that scale alone does not guarantee alignment with moral attitudes if the underlying training data or methodology is insufficiently diverse or biased.
@@
In addition, topic-level analysis reveals that certain issues (e.g., political violence, terrorism, or wife-beating) consistently produce higher mean errors across different architectures. These discrepancies suggest that moral questions involving violence or extreme social norms may pose particular challenges for current language models, especially when training data do not include nuanced representations of such topics. Even models that perform relatively well on broad measures sometimes fail on region-specific or contentious issues. This trend aligns with evidence that LLMs handle clear-cut moral scenarios well but often display uncertainty or divergence on morally ambiguous dilemmas (Scherrer et al., 2023). Per-country heatmaps similarly highlight that no single model excels in all areas: while a model may align with opinions in Western nations, it can deviate markedly in communities whose moral or cultural practices are underrepresented in its training corpora.
@@
Despite these limitations, instruction-tuned and larger models show promise in better reflecting overall moral consensus in many cases. This suggests that scaling models and using tailored training, where instructions or datasets capture diverse viewpoints, can improve moral judgment alignment. However, performance still varies, highlighting the need to analyze results in detail (e.g., by topic or country) rather than relying on a single global metric. From an applied perspective, these insights can guide the development of more culturally responsive AI systems, for example, informing content moderation policies or chatbot designs that respect regional norms.”
5.1 Limitations
Although our methodology offers insights into cross-cultural moral alignment in language models, it has several limitations that should be acknowledged. First, the WVS and PEW data capture broad national averages and may not fully reflect within-country heterogeneity, especially in regions with significant cultural or linguistic diversity. Second, our log-probability difference calculation relies on short prompt templates, which might not elicit the full context required for more complex moral issues. Third, the models we evaluated differ in size, instruction tuning, and training data composition, making it challenging to isolate the effect of each factor.
A further limitation arises from the necessity of employing distinct evaluation strategies. For local models, we have access to token-level log probabilities, enabling us to compute log-probability differences as a proxy for moral judgment. However, for OpenAI’s proprietary chat models, we rely on directly elicited numerical scores because the API does not expose internal log probabilities. This divergence means that the resulting moral scores are derived from different underlying mechanisms, precluding a direct, unified comparison of model outputs in our visualizations. Future work might seek alternative methods to bridge this gap or develop metrics that are comparable across elicitation approaches.
6 Conclusion
@@
In conclusion, our analysis of moral stance alignment across WVS and PEW data underscores both the progress and the continuing gaps in LLMs’ performance. Models with substantial parameter counts and instruction-tuned frameworks frequently achieve moderate-to-high correlations with surveyed human judgments, suggesting an ability to capture broad moral viewpoints. However, sizable deviations persist on sensitive topics and in particular cultural contexts, indicating that no current model entirely overcomes biases or data deficiencies. Thus, while larger or more specialized training procedures can improve a model’s capacity to reflect human moral attitudes, they do not guarantee universal alignment. Future work must address these persistent shortcomings through expanded training corpora, targeted bias mitigation, and refined evaluation protocols that account for cultural and topic-level nuances.
Ethical considerations
Using language models in real-world applications has important ethical implications and risks. Even though these models can approximate broad moral opinions, they may misrepresent local or minority viewpoints if their training data is not diverse enough. This misrepresentation can lead to biases or stereotypes, especially on sensitive topics like domestic violence, religious norms, or political extremism. If a model’s output is mistakenly viewed as a true reflection of public opinion, automated decisions could unfairly target or exclude certain groups, worsening existing inequalities. Moreover, significant misalignment on controversial topics can undermine public trust if model predictions seem harmful or insensitive. To reduce such risks, it is vital to include diverse voices and expert feedback when building and testing these models. Adding regular evaluations on moral or cultural issues, transparent reports of known biases, and human review for high-stakes decisions, can help ensure ethical and responsible deployment. As language models evolve, balancing technical progress with careful oversight will be essential for maintaining fairness and trust in automated systems.
Funding
This research received no external funding.
Disclosure statement
The authors declare no conflict of interest.
Data and code availability
The full source code, experiment scripts, and processed datasets are openly available on GitHub. [^4] https://github.com/mohammadi-hadi/cultural-moral-judgments-llms
Author Contributions
H.M. and A.B. conceptualized the research. H.M., E.P., Y.M., and A.B. developed the methodology. H.M., E.P., and Y.M. contributed to software implementation, while H.M. and A.B. handled validation. E.P., Y.M., and H.M. composed the original draft, and H.M. and A.B. oversaw review and editing.
@@ Acknowledgments
We appreciate the maintainers of WVS and PEW data for enabling large-scale cross-cultural analysis. We also thank Dr. Anastasia Giachanou for her valuable comments and feedback and Tina Shahedi for her editorial contributions to this paper. We also thank SURF for providing the computational facilities.
Appendix A Topic Codes for WVS and PEW
Table 4: Mapping of Topic Codes to the Dataset (WVS or PEW) and their corresponding moral questions.
Topic Code Dataset Moral Question
Q177 WVS Claiming government benefits to which you are not entitled
Q178 WVS Avoiding a fare on public transport
Q179 WVS Stealing property
Q180 WVS Cheating on taxes
Q181 WVS Someone accepting a bribe in the course of their duties
Q182 WVS Homosexuality
Q183 WVS Prostitution
Q184 WVS Abortion
Q185 WVS Divorce
Q186 WVS Sex before marriage
Q187 WVS Suicide
Q188 WVS Euthanasia
Q189 WVS For a man to beat his wife
Q190 WVS Parents beating children
Q191 WVS Violence against other people
Q192 WVS Terrorism as a political, ideological or religious mean
Q193 WVS Having casual sex
Q194 WVS Political violence
Q195 WVS Death penalty
Q84A PEW Using contraceptives
Q84B PEW Getting a divorce
Q84C PEW Having an abortion
Q84D PEW Homosexuality
Q84E PEW Drinking alcohol
Q84F PEW Married people having an affair
Q84G PEW Gambling
Q84H PEW Sex between unmarried adults
Appendix B WVS & PEW scores by country
Figure 8 compares normalized WVS (orange) and PEW (gold) scores by country. Each box shows the interquartile range, with medians as horizontal lines and diamonds marking outliers. The broader spread in the WVS data for many countries suggests higher variance in moral acceptance. Some countries, such as the United States or Czech Republic, show very wide ranges, from near $-1$ (never justifiable) to close to $+1$ (always justifiable). Others, often in the Middle East or South Asia, have more negative medians, reflecting stricter cultural norms on certain issues.
Figure 8: Distribution of normalized WVS (orange) and PEW (gold) survey scores by country.
Appendix C Individual Figures by Model & Dataset
In each scatter plot, the horizontal axis survey_score corresponds to WVS in Figure 9 and PEW ratings in Figure 10. Meanwhile, the vertical axis log_prob_diff shows the difference between the log-probability the model assigns to a morally justifiable statement vs. a morally unjustifiable statement. A positive slope suggests that higher survey acceptance correlates with higher log-prob differences in the same direction, meaning better alignment. Conversely, negative slopes may show systematic misalignment on that dimension.
GPT2-B
GPT2-M
GPT2-L
OPT-125
OPT-350
BloomZ
Qwen-0.5
Qwen-72
Llama3-8B
Llama3.3-70I
Llama3.3-70I
Falcon3-7B
Falcon-40I
GPT-NeoX20
Dolly-12
Bloom
Llama2-70
Figure 9: Scatter plots for WVS dataset
GPT2-B
GPT2-M
GPT2-L
OPT-125
OPT-350
BloomZ
Qwen-0.5
Qwen-72
Llama3-8B
Llama3.3-70I
Llama3.3-70I
Falcon3-7B
Falcon-40I
GPT-NeoX20
Dolly-12
Bloom
Llama2-70
Figure 10: Scatter plots for PEW dataset

### v2

@@ 5 Discussion
These results reveal a complex picture that merits careful interpretation. Our findings show considerable variation in how well language models replicate cross-cultural moral judgments, as captured in the WVS and PEW surveys. Larger or instruction-tuned models, such as Falcon-40B-Inst, Gemma-2-9B-IT, and GPT-4o, frequently demonstrate higher correlations with aggregated human survey responses. In contrast, some models, including Qwen-0.5B and Llama-2-70B, yield systematically negative correlations, suggesting that scale alone does not guarantee alignment with moral attitudes if the underlying training data or methodology is insufficiently diverse or biased.
@@
Topic-level analysis reveals that certain issues, such as political violence, terrorism, or wife-beating, consistently produce higher mean errors across different architectures. These discrepancies suggest that moral questions involving violence or extreme social norms may pose particular challenges for current language models, especially when training data do not include nuanced representations of such topics. Even models that perform relatively well on broad measures sometimes fail on region-specific or contentious issues. This trend aligns with evidence that LLMs handle clear-cut moral scenarios well but often display uncertainty or divergence on morally ambiguous dilemmas (?). Recent work by ? examines whether LLMs possess intrinsic self-correction capabilities for moral reasoning, finding that such mechanisms are often superficial rather than genuinely reflecting moral understanding.
We can offer a tentative explanation for why certain topics are harder than others. Topics like political violence, terrorism, and suicide are rarely discussed approvingly in published text, so models learn near-absolute condemnation and miss cultural nuances about resistance movements or end-of-life decisions. Wife-beating is universally condemned in formal text, yet cultural practices vary; models cannot bridge this gap between published norms and lived reality. In contrast, topics like divorce, alcohol consumption, and contraceptive use are widely discussed with clear cultural variation, allowing models to learn regional differences from their training data.
Our regional analysis confirms a substantial W.E.I.R.D. bias in current LLMs. Models align best with Western European and North American perspectives, while Sub-Saharan African and MENA regions show the weakest alignment. This finding has important implications for global deployment of LLM-based systems: users from underrepresented regions may receive responses that do not reflect their cultural values or may even contradict local moral norms.
@@
Despite these limitations, instruction-tuned and larger models show promise in better reflecting moral consensus in many cases. This suggests that scaling models and using tailored training that captures diverse viewpoints can improve moral judgment alignment. However, performance still varies, highlighting the need to analyze results in detail (e.g., by topic or country) rather than relying on a single global metric. From an applied perspective, these insights can guide the development of more culturally responsive AI systems, for example, informing content moderation policies or chatbot designs that respect regional norms.
For practitioners deploying LLMs in global contexts, our findings suggest three actionable recommendations: (1) implement region-specific calibration for morally sensitive applications, rather than assuming a single model configuration works universally; (2) consider ensemble approaches that combine predictions from models trained on different cultural corpora, particularly for underrepresented regions; and (3) establish human-in-the-loop validation for high-stakes moral judgments, especially when serving users from Sub-Saharan Africa, MENA, or other regions where model alignment is weakest.
6 Conclusion and Limitations
@@
Our analysis of moral stance alignment across WVS and PEW data underscores both the progress and the continuing gaps in LLMs’ performance. Models with substantial parameter counts and instruction-tuned frameworks frequently achieve moderate-to-high correlations with surveyed human judgments, suggesting an ability to capture broad moral viewpoints. However, sizable deviations persist on sensitive topics and in particular cultural contexts, indicating that no current model entirely overcomes biases or data deficiencies. Thus, while larger or more specialized training procedures can improve a model’s capacity to reflect human moral attitudes, they do not guarantee universal alignment. Future work must address these persistent shortcomings through expanded training corpora, targeted bias mitigation, and refined evaluation protocols that account for cultural and topic-level nuances.
In summary, current LLMs are not culturally neutral arbiters of morality—they reflect the values embedded in their predominantly Western training data. Until training pipelines achieve genuine cultural diversity, applications involving moral judgments should be deployed with explicit regional validation and clear user awareness of potential cultural biases.
Limitations.
Although our methodology offers insights into cross-cultural moral alignment in language models, it has several limitations that should be acknowledged. First, the WVS and PEW data capture broad national averages and may not fully reflect within-country heterogeneity, especially in regions with significant cultural or linguistic diversity. Second, our log-probability difference calculation relies on short prompt templates, which might not elicit the full context required for more complex moral issues. Future work could explore richer contextual prompts or narrative-based evaluation methods. Third, the models we evaluated differ in size, instruction tuning, and training data composition, making it challenging to isolate the effect of each factor. Fourth, the min-max normalization we apply does not fully address cultural differences in scale usage. Fifth, as noted earlier, potential data leakage from survey reports into model training data may inflate alignment scores, particularly for proprietary models. Future work could establish standardized metrics that enable more consistent comparisons across models and prompting approaches.
@@ Acknowledgments
We thank Efthymia Papadopoulou and Yasmeen F.S.S. Meijer for their contributions to the exploratory data analysis and earlier analysis. We thank the maintainers of the WVS and PEW data for enabling large-scale cross-cultural analysis. We also thank anonymous reviewers for their valuable comments and feedback. Computational resources were provided by SURF.